OpenStudy (sbuck98):

anyone want to help with 5 polynomials for medals &coins?

OpenStudy (yuliacosta111):

me

OpenStudy (sbuck98):

@Yuliacosta111

OpenStudy (sbuck98):

@imqwerty

OpenStudy (mathmale):

These algebra problems hinge on the Distributive Property of Multiplication. Perhaps you've heard of the "FOIL" method (first, outer, inner, last). Example: (a+b)(c+d)=ac+ad+bc+bd. Have you? Have you any examples to refer to in your class materials?

OpenStudy (sbuck98):

No:( @mathmale

OpenStudy (mathmale):

Happy to help...but I do expect that you do some basic reading and look for examples of algebraic multiplication (including the FOIL method) beforehand. Willing to do that?

OpenStudy (mathmale):

Please see the following, online: http://www.regentsprep.org/regents/math/algebra/av3/smul_bin.htm

OpenStudy (mathmale):

Example 2: (x+2y)(x-y)=x^2+2xy-xy-2y^2), using the same FOIL method. The First terms are x and x; their product is x^2. The Outer terms are x and -y; their product is -xy. The Inner terms are 2y and x; their product is 2xy. The Last terms are 2y and -y; their product is -2y^2. Just combine these four products to obtain the final answer. Apply this procedure to the problems you've posted. I apologize, but will have to log off the 'Net now. Good luck!